alis:

A spinner has 8 equal sectors labeled from 1 to 8. The spinner is spun twice. What is the probability of getting an even number on the first spin and another even number on the second spin?

Vocaloid:

for this event you need to consider both spins separately what's the probability of getting even on one spinner with sections labelled 1-8?

RawlsK29:

Wouldn't you have like a 4/8 chance of landing on an even number

Vocaloid:

let's let the op try to figure it out ^_^

alis:

sorry i was doing something but umm i'd be guessing. would it be an odd number out of 8?

Vocaloid:

let's list out the outcomes for the spinner: 1,2,3,4,5,6,7,8 how many even numbers are there?

alis:

4

alis:

oh so it'd be 4 sorry i'm dumb lol

Vocaloid:

good, so for the first spinner the probability is (4/8) or (1/2) for the second spinner, the probability is the same (independent events) so (1/2)*(1/2) = (1/4) = your answer

RawlsK29:

Thats what i was thinking, but it sounded to easy at first.

alis:

omg that's actually very simple thank you so much